# On Solving Asymmetric Diagonally Dominant Linear Systems in Sublinear Time

### v1

@@ 1 Introduction
In particular, for the important classes of Laplacian and symmetric diagonally dominant (SDD) systems, the breakthrough work of Spielman and Teng [ST04, ST14] established the first nearly-linear-time (in $\operatorname{nnz}(\mathbf{M})$) solvers. This gave rise to the influential Laplacian Paradigm, which revolutionized algorithmic graph theory and numerical linear algebra, with widespread applications ranging from network science to machine learning (see, e.g., [Ten10, Vis13]). Among subsequent efforts to generalize the SDD solvers, a line of work [CKP+16, CKP+17, CKK+18] developed nearly-linear-time solvers for asymmetric row/column diagonally dominant (RDD/CDD) systems, which significantly expanded the scope of the Laplacian Paradigm.
@@
Our algorithmic toolkit and investigation of the diverse upper bounds are inspired by recent advances in local algorithms for estimating PageRank [BP98] and effective resistance [DS84] on graphs [YWW+24, WWY24, WWWY24, Wan24, CWW25, YLLW25], which are important special cases of solving RDD/CDD systems. The techniques of random-walk sampling [FRCS05], local push [ACL07, ABC+08], and the bidirectional method [LBG16] have been extensively studied for these problems, and recent works have further uncovered their new properties and optimality in certain settings. Nonetheless, previous works typically analyze their applications to PageRank and effective resistance computation separately. Our perspective of formulating these problems as solving linear systems, however, provides a more general and unified framework for understanding these techniques and problems, revealing their deeper connections. As we shall see, this bigger picture yields novel insights, extended results, and improved complexity bounds for existing problems.
Notably, our perspective reveals a connection between two fundamental local push algorithms on graphs, namely ForwardPush [ACL07] and BackwardPush [ABC+08]. These algorithms iteratively perform local push operations to explore the graph in opposite directions. Although both algorithms share similar approaches, they have been treated as distinct methods for different problems, with each analyzed separately. In contrast, by abstracting both methods as a single algebraic primitive, we demonstrate that ForwardPush and BackwardPush are equivalent to applying this primitive to different linear systems. This characterization helps to explain their distinct properties and enables unified analysis of both approaches within our framework.
@@
In the remainder of this section, we formally define the problem and present our main contributions, followed by a technical overview and discussion of future directions.
@@ 1.2 Problem Formulation
For effective resistance, we consider a connected undirected graph $G$ and two distinct nodes $s,t\in V$. The effective resistance (a.k.a. resistance distance) between $s$ and $t$, denoted by $R_{G}(s,t)$, is defined as the equivalent resistance between $s,t$ if the graph is thought of as an electrical network with edge $(u,v)\in E$ having resistance $1/\mathbf{A}_{G}(u,v)$. Algebraically, $R_{G}(s,t)=(\boldsymbol{e}_{s}-\boldsymbol{e}_{t})^{\top}\mathbf{L}_{G}^{+}(\boldsymbol{e}_{s}-\boldsymbol{e}_{t})$. As we will establish in Lemma 9.1, setting $\mathbf{M}=\mathbf{L}_{G}$ and $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$ in our formulation yields $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}=R_{G}(s,t)$. Here, $\mathbf{M}=\mathbf{L}_{G}$ is SDDZ and $\boldsymbol{b}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}\in\operatorname{range}(\mathbf{M})$.
@@ 1.3 Previous Work for SDD Systems
On the negative side, [AKP19] proves an $\Omega\left(\kappa(\mathbf{S})^{2}/\log^{3}n\right)=\widetilde{\Omega}\left(\kappa(\mathbf{S})^{2}\right)$ query lower bound (in terms of probing $\boldsymbol{b}$) for achieving a weaker absolute error bound of $\varepsilon\left\|\boldsymbol{x}^{\ast}\right\|_{\infty}$, for $\kappa(\mathbf{S})=O(\sqrt{n}/\log n)$ and $\varepsilon=\Theta(1/\log n)$. The matrix $\mathbf{S}$ in their hard instance is a Laplacian matrix of a fixed undirected unweighted graph with maximum degree $4$ and thus satisfies $\kappa(\widetilde{\mathbf{S}})=\Theta\big{(}\kappa(\mathbf{S})\big{)}$. Therefore, this lower bound can also be written as $\widetilde{\Omega}\left(1/\gamma(\mathbf{S})^{2}\right)$. To our knowledge, no other work has explicitly studied sublinear-time SDD/RDD/CDD solvers.
1.4 Formulation of $\boldsymbol{x}^{\ast}$ and the $p$-Norm Gap
@@
where $q$ is conjugate to $p$. We further define the maximum $p$-norm gap of $\mathbf{M}$ as $\gamma_{\max}(\mathbf{M}):=\max_{p\in[1,\infty]}\gamma_{p}(\mathbf{M})$. [^2] Alternatively, one can define $\gamma_{\max}(\mathbf{M}):=\max_{p\in\{1,2,\infty\}}\gamma_{p}(\mathbf{M})$, which keeps our results intact. To our knowledge, no prior work has explicitly studied these quantities.
@@ 1.5 Main Algorithmic Results
This theorem subsumes the main algorithmic result of [AKP19] for SDD systems and extends it to the RDD case while achieving a $(\log L)$-factor improvement. Their original result is recovered as a special case when $\mathbf{M}$ is SDD and $\boldsymbol{t}$ is a canonical unit vector. On the other hand, the lower bound in [AKP19] implies that a quadratic dependence on $1/\gamma$ in our upper bound is necessary.
We also show that for RDDZ $\mathbf{M}$ along with nonnegative vectors $\boldsymbol{b}$ and $\boldsymbol{t}$, if we allow the relaxed error bound $\varepsilon\|\boldsymbol{x}^{\ast}\|_{\infty}$, the complexity can be improved to depend quadratically on $L$. We adopt this error measure since it provides a natural accuracy guarantee and has been previously studied in [AKP19]. Lemma 5.2 will establish that $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\leq 2\|\boldsymbol{x}^{\ast}\|_{\infty}$ for RDD systems.
@@
The following theorem further considers the relative error guarantee of $\varepsilon\cdot\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$. The complexity still depends quadratically on $L$ and it also depends on $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\big{/}\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$.
@@
Lastly, applying the bidirectional method to special RCDD systems yields complexity bounds with improved dependence on $L$ and $\varepsilon$, achieving either $L^{7/3}\varepsilon^{-2/3}$ or $L^{5/2}\varepsilon^{-1}$ factors.
@@
As we shall see, all theorems in this subsection except Theorem 1.5 hold if we replace RDD/RDDZ by CDD/CDDZ, swap $\boldsymbol{b}$ and $\boldsymbol{t}$ (except in $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$), and replace $f_{\mathrm{row}}(\mathbf{M})$ by $f_{\mathrm{col}}(\mathbf{M})$ in the statements. Theorem 1.5 is an exception because its proof relies on the property that $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\leq 2\|\boldsymbol{x}^{\ast}\|_{\infty}$ for RDD $\mathbf{M}$, which does not have a straightforward analog for the CDD case.
@@ 1.6 Connections to PageRank and Effective Resistance Computation
Our results can be directly applied to the PPR or PageRank contribution equations and the effective resistance problem to yield complexity bounds for different graph types under various access models and accuracy guarantees. Rather than exhaustively presenting all applications, here we highlight selected results that generalize and improve upon the previously best complexity bounds. Additional applications and comparisons are provided in Sections 8 and 9.
@@
This improves over the previously best upper bounds of $O\left(\frac{1}{\varepsilon^{2}\delta_{G}}\cdot\min\left(\frac{d_{G}(t)}{\alpha^{2}},\frac{\sqrt{m}}{\alpha^{2}}\right)\right)$, which was given by [Wan24] and stated for unweighted undirected graphs.
For effective resistance computation, recall that we consider connected undirected graphs $G$. Lemma 9.1 establishes that the value $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$ that our algorithms approximate equals $R_{G}(s,t)$ when setting $\mathbf{M}=\mathbf{L}_{G}$ and $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$. Thus, Theorems 1.4, 1.7, and 1.8 directly imply the following complexity bounds for estimating effective resistance.
@@
For any connected undirected unweighted graph $G$, given nodes $s,t\in V$ and $\gamma>0$ as a lower bound on $\gamma(\mathbf{L}_{G})$, there exists a randomized algorithm that estimates the effective resistance $R_{G}(s,t)$ within absolute error $\varepsilon$ with probability at least $3/4$ in time
$O\left(\min\left(\frac{L^{3}}{\varepsilon^{2}\cdot\min\big{(}d_{G}(s),d_{G}(t)\big{)}^{2}},\frac{L^{3}}{\varepsilon},\frac{L^{7/3}}{\varepsilon^{2/3}},\frac{L^{5/2}}{\varepsilon\cdot\min\big{(}d_{G}(s),d_{G}(t)\big{)}^{1/2}}\right)\right),$
@@
This result subsumes the previously best upper bounds of $O\left(\min\left(\frac{L^{3}\log L}{\varepsilon^{2}\cdot\min(d_{G}(s),d_{G}(t))^{2}},\frac{L^{7/3}\log L}{\varepsilon^{2/3}}\right)\right)$ given by [CWW25] with essentially the same setting of $L$. Moreover, since $R_{G}(s,t)$ can be lower bounded by $1/2\big{/}\min\big{(}d_{G}(s),d_{G}(t)\big{)}$ [Lov93, Corollary 3.3], by setting the absolute error parameter $\varepsilon$ to be $\varepsilon_{r}\cdot 1/2\big{/}\min\big{(}d_{G}(s),d_{G}(t)\big{)}$, the last bound in our result implies that an estimate of $R_{G}(s,t)$ within relative error $\varepsilon_{r}$ can be computed in time $O\left(\min\big{(}d_{G}(s),d_{G}(t)\big{)}^{1/2}L^{5/2}\varepsilon_{r}^{-1}\right)$. This improves over the previously best upper bound of $O\left(\min\big{(}d_{G}(s),d_{G}(t)\big{)}^{1/2}L^{3}\log L\cdot\varepsilon_{r}^{-1}\right)$ given by [YLLW25].
@@
For any large enough $n$ and $\varepsilon=\Omega(1/n)$, there exist $\boldsymbol{b}\in\mathbb{R}^{n}$ and $\boldsymbol{t}\in[n]$ that satisfy the following. Every randomized algorithm that, given access to an invertible SDD matrix $\mathbf{S}\in\mathbb{R}^{n\times n}$ whose spectral gap is $\Omega(1)$, succeeds with probability at least $3/4$ to approximate $\boldsymbol{x}^{\ast}(t)$ within absolute error $\varepsilon\|\boldsymbol{x}^{\ast}\|_{\infty}$, must probe $\Omega(1/\varepsilon)$ coordinates of $\mathbf{S}$ in the worst case. Here, $\boldsymbol{x}^{\ast}=\mathbf{S}^{-1}\boldsymbol{b}$.
@@ 1.8 Technical Overview
Our algorithms estimate $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}=\frac{1}{2}\boldsymbol{t}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ by adapting three techniques for random-walk probability estimation on graphs: random-walk sampling [FRCS05], local push [ACL07, ABC+08], and the bidirectional method [LBG16].
@@
Based on local push methods, we formulate a Push primitive that estimates the vector $\boldsymbol{x}^{\ast}_{L}=\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ through deterministic local computation. Push maintains coordinate variables initialized to $\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ and iteratively applies the linear operator $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)$ to selected coordinates. Our analysis relies on invariant properties preserved by these operations, including an inequality variant that handles negative entries. This approach yields closed-form accuracy guarantees for RDD systems and complexity bounds for special CDD systems.
The bidirectional method combines random-walk sampling and local push from two directions. We adapt the BiPPR framework [LBG16], using the invariant property of Push from $\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ to construct an unbiased estimator for $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$, which can be sampled via random walks from $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$ when $\mathbf{M}$ is RDD. Proper parameter selection balances the computational costs of both components, achieving improved dependence on $L$ and $\varepsilon$, particularly for RCDD systems.

### v2

@@ 1 Introduction
In particular, for the important classes of Laplacian and symmetric diagonally dominant (SDD) systems, the breakthrough work of Spielman and Teng [ST04, ST14] established the first nearly-linear-time (in $\operatorname{nnz}(\mathbf{M})$) solvers. This gave rise to the influential Laplacian Paradigm, which revolutionized algorithmic graph theory and numerical linear algebra, with widespread applications ranging from network science to machine learning (see, e.g., [Ten10, Vis13]). Among subsequent efforts to generalize the SDD solvers, a line of work [CKPPSV16, CKPPRSV17, CKKPPRS18] developed nearly-linear-time solvers for asymmetric row/column diagonally dominant (RDD/CDD) systems, which significantly expanded the scope of the Laplacian Paradigm.
@@
Our algorithmic toolkit and investigation of the diverse upper bounds are inspired by recent advances in local algorithms for estimating PageRank [BP98] and effective resistance [DS84] on graphs [YWWWW24, WWY24, WWWY24, Wan24, BJTWY25, CWW25, YLLW25, TWWY26], which are important special cases of solving RDD/CDD systems. The techniques of random-walk sampling [ST04, FRCS05], local push [ACL07, ABCHMT08], and the bidirectional method [LBG16] have been extensively studied for these problems, and these recent works have further uncovered their new properties and optimality in certain settings. Nonetheless, previous works typically analyze their applications to PageRank and effective resistance separately. Our perspective of formulating these problems as solving linear systems, however, provides a more general and unified framework for understanding these techniques and problems, revealing their deeper connections. As we shall see, this bigger picture yields novel insights, extended results, and improved complexity bounds.
Notably, our perspective reveals a connection between two fundamental local push algorithms on graphs, namely ForwardPush [ACL07] and BackwardPush [ABCHMT08]. These algorithms iteratively perform local push operations to explore the graph in opposite directions. Although both algorithms share similar approaches, they have been treated as distinct methods for different problems, with each analyzed separately. In contrast, by abstracting both methods as a single algebraic primitive, we demonstrate that ForwardPush and BackwardPush are equivalent to applying this primitive to different linear systems. This characterization helps to explain their distinct properties and enables unified analysis of both approaches.
@@
In the remainder of this section, we formally define the problem, present our main contributions, and provide a technical overview and discussion of future directions.
@@ 1.2 Problem Formulation
For effective resistance, we consider a connected undirected graph $G$ and two distinct nodes $s,t\in V$. The effective resistance (a.k.a. resistance distance) between $s$ and $t$, denoted by $R_{G}(s,t)$, is defined as the equivalent resistance between $s,t$ if the graph is thought of as an electrical network with each edge $(u,v)\in E$ having resistance $1/\mathbf{A}_{G}(u,v)$. Algebraically, $R_{G}(s,t)=(\boldsymbol{e}_{s}-\boldsymbol{e}_{t})^{\top}\mathbf{L}_{G}^{+}(\boldsymbol{e}_{s}-\boldsymbol{e}_{t})$. As we will establish in Lemma 9.1, setting $\mathbf{M}=\mathbf{L}_{G}$ and $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$ in our formulation yields $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}=R_{G}(s,t)$. Here, $\mathbf{M}=\mathbf{L}_{G}$ is SDDZ and $\boldsymbol{b}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}\in\operatorname{range}(\mathbf{M})$.
@@ 1.3 Previous Work for SDD Systems
On the negative side, [AKP19] proves an $\Omega\left(\kappa(\mathbf{S})^{2}/\log^{3}n\right)=\widetilde{\Omega}\left(\kappa(\mathbf{S})^{2}\right)$ query lower bound (in terms of probing $\boldsymbol{b}$) for achieving a weaker absolute error bound of $\varepsilon\left\|\boldsymbol{x}^{\ast}\right\|_{\infty}$, for $\kappa(\mathbf{S})=O(\sqrt{n}/\log n)$ and $\varepsilon=\Theta(1/\log n)$. The matrix $\mathbf{S}$ in their hard instance is a Laplacian matrix of a fixed unweighted undirected graph with maximum degree $4$ and thus satisfies $\kappa(\widetilde{\mathbf{S}})=\Theta\big(\kappa(\mathbf{S})\big)$. Therefore, this lower bound can also be written as $\widetilde{\Omega}\left(1/\gamma(\mathbf{S})^{2}\right)$. To our knowledge, no other work has explicitly studied sublinear-time SDD/RDD/CDD solvers.
1.4 Formulation of $\boldsymbol{x}^{\ast}$ and the $p$-Norm Gaps
@@
where $q$ is conjugate to $p$. We further define the maximum $p$-norm gap of $\mathbf{M}$ as $\gamma_{\max}(\mathbf{M}):=\max_{p\in[1,\infty]}\gamma_{p}(\mathbf{M})$. [^2] Alternatively, one can define $\gamma_{\max}(\mathbf{M}):=\max_{p\in\{1,2,\infty\}}\gamma_{p}(\mathbf{M})$, which yields a possibly smaller quantity but our main results continue to hold with this definition. To our knowledge, no prior work has explicitly studied these quantities.
@@ 1.5 Main Algorithmic Results
This theorem subsumes the main algorithmic result of [AKP19] for SDD systems and extends it to the RDD case while achieving a mild $(\log L)$-factor improvement. Their original result is recovered as a special case when $\mathbf{M}$ is SDD and $\boldsymbol{t}$ is a canonical unit vector. On the other hand, the lower bound in [AKP19] implies that a quadratic dependence on $1/\gamma$ in the compelxity bound is necessary. Our algorithm is similar to theirs, but we achieve the improved complexity by adopting a different random-walk sampling scheme.
We also show that for RDDZ $\mathbf{M}$ along with nonnegative vectors $\boldsymbol{b}$ and $\boldsymbol{t}$, if we allow the relaxed error bound $\varepsilon\|\boldsymbol{x}^{\ast}\|_{\infty}$, the complexity can be improved to depend quadratically on $L$ via a variance analysis of the random-walk sampling process. We adopt this error measure since it provides a natural accuracy guarantee and has been previously studied in [AKP19]. Lemma 5.2 will establish that $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\leq 2\|\boldsymbol{x}^{\ast}\|_{\infty}$ for RDD systems.
@@
The following theorem further considers the relative error guarantee of $\varepsilon\cdot\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$. The complexity depends quadratically on $L$ and linearly on $\|\boldsymbol{t}\|_{1}\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\big/\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$ and is also achieved using random-walk sampling.
@@
Lastly, applying the bidirectional method to special RCDD systems yields complexity bounds with improved dependence on $L$ and $\varepsilon$, achieving either $L^{7/3}\varepsilon^{-2/3}$ or $L^{5/2}\varepsilon^{-1}$ using different parameter settings.
@@
As we shall see, all theorems in this subsection except Theorem 1.5 still hold if we replace RDD/RDDZ by CDD/CDDZ, swap $\boldsymbol{b}$ and $\boldsymbol{t}$ (except in $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$), and replace $f_{\mathrm{row}}(\mathbf{M})$ by $f_{\mathrm{col}}(\mathbf{M})$ in the statements. Theorem 1.5 is an exception because its proof relies on the property that $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\leq 2\|\boldsymbol{x}^{\ast}\|_{\infty}$ for RDD $\mathbf{M}$, which does not have a simple analog for the CDD case.
@@ 1.6 Connections to PageRank and Effective Resistance Computation
Our results can be directly applied to the PPR or PageRank contribution equations and the single-pair effective resistance problem to yield complexity bounds for different graph types under various access models and accuracy guarantees. Notably, as we shall see, for PageRank computation, half the decay factor serves as a lower bound on the maximum $p$-norm gap of the corresponding systems; for effective resistance computation, the spectral gap of the graph Laplacian equals the maximum $p$-norm gap of the system. Rather than exhaustively presenting all applications of our results, here we highlight selected results that generalize and improve upon the previously best complexity bounds. Additional applications and comparisons are provided in Sections 8 and 9.
@@
This improves over the previously best upper bounds of $O\left(\frac{1}{\varepsilon^{2}\delta_{G}}\cdot\min\left(\frac{d_{G}(t)}{\alpha^{2}},\frac{\sqrt{m}}{\alpha^{2}}\right)\right)$, which was given by [Wan24] and stated for unweighted undirected graphs. Our algorithm is essentially the same as that in [Wan24], which generates random walks from the target node $t$, and the improvement comes from our tighter lower bounds on $\boldsymbol{\pi}_{G,\alpha}(t)$ for Eulerian graphs.
For effective resistance computation, recall that we consider connected undirected graphs $G$. Lemma 9.1 establishes that the value $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}$ that our algorithms approximate equals $R_{G}(s,t)$ when setting $\mathbf{M}=\mathbf{L}_{G}$ and $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$. Thus, Theorems 1.4 and 1.8 directly imply the following complexity bounds for estimating effective resistance.
@@
For any connected unweighted undirected graph $G$, given nodes $s,t\in V$ and $\gamma>0$ as a lower bound on $\gamma(\mathbf{L}_{G})$, there exists a randomized algorithm that estimates the effective resistance $R_{G}(s,t)$ within absolute error $\varepsilon$ with probability at least $3/4$ in time
$O\left(\min\left(\frac{L^{3}}{\varepsilon^{2}\cdot\min\big(d_{G}(s),d_{G}(t)\big)^{2}},\frac{L^{7/3}}{\varepsilon^{2/3}},\frac{L^{5/2}}{\varepsilon\cdot\min\big(d_{G}(s),d_{G}(t)\big)^{1/2}}\right)\right),$
@@
This result subsumes the previously best upper bounds of $O\left(\min\left(\frac{L^{3}\log L}{\varepsilon^{2}\cdot\min(d_{G}(s),d_{G}(t))^{2}},\frac{L^{7/3}\log L}{\varepsilon^{2/3}}\right)\right)$ given by [CWW25] with essentially the same setting of $L$. Moreover, since $R_{G}(s,t)$ can be lower bounded by $1/2\big/\min\big(d_{G}(s),d_{G}(t)\big)$ [Lov93, Corollary 3.3], by setting the absolute error parameter $\varepsilon$ to be $\varepsilon_{r}\cdot 1/2\big/\min\big(d_{G}(s),d_{G}(t)\big)$, the last bound in our result implies that an estimate of $R_{G}(s,t)$ within relative error $\varepsilon_{r}$ can be computed in time $O\left(\min\big(d_{G}(s),d_{G}(t)\big)^{1/2}L^{5/2}\varepsilon_{r}^{-1}\right)$. This improves over the previously best upper bound of $O\left(\min\big(d_{G}(s),d_{G}(t)\big)^{1/2}L^{3}\log L\cdot\varepsilon_{r}^{-1}\right)$ given by [YLLW25]. Our algorithms are essentially the same as those in [CWW25, YLLW25], i.e., using random-walk sampling and the bidirectional method, and the improvements stem from our simpler sampling scheme that avoids using extra data structures and a refined analysis.
@@
For any large enough $n$ and $\varepsilon=\Omega(1/n)$, there exist $\boldsymbol{b}\in\mathbb{R}^{n}$ and $t\in[n]$ that satisfy the following. Every randomized algorithm that, given access to an invertible SDD matrix $\mathbf{S}\in\mathbb{R}^{n\times n}$ whose spectral gap is $\Omega(1)$, succeeds with probability at least $3/4$ to approximate $\boldsymbol{x}^{\ast}(t)$ within absolute error $\varepsilon\|\boldsymbol{x}^{\ast}\|_{\infty}$, must probe $\Omega(1/\varepsilon)$ coordinates of $\mathbf{S}$ in the worst case. Here, $\boldsymbol{x}^{\ast}=\mathbf{S}^{-1}\boldsymbol{b}$.
@@ 1.8 Technical Overview
Specifically, by the decomposition $\mathbf{M}=\mathbf{D}_{\mathbf{M}}-\mathbf{A}_{\mathbf{M}}^{\top}$, the equation $\mathbf{M}\boldsymbol{x}=\boldsymbol{b}$ is equivalent to $\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{M}\boldsymbol{x}=\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$, i.e., $\left(\mathbf{I}-\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\boldsymbol{x}=\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$. For certain types of $\mathbf{M}$, $\mathbf{I}-\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}$ is invertible and we can write the unique solution $\boldsymbol{x}^{\ast}$ as $\boldsymbol{x}^{\ast}=\left(\mathbf{I}-\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)^{-1}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}=\sum_{\ell=0}^{\infty}\left(\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$. Our formulation of the specific solution $\boldsymbol{x}^{\ast}=\frac{1}{2}\sum_{\ell=0}^{\infty}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ in Theorem 1.1 is similar to this expression but uses the “lazified” matrix $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)$ instead of $\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}$, which helps to ensure convergence for general RDD/CDD $\mathbf{M}$.
Furthermore, to analyze the truncation error, intuitively, we require a quantity that measures the strength of the diagonal dominance property of $\mathbf{M}$. If $\mathbf{M}$ is strictly RDD, a natural choice is $1-\left\|\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right\|_{\infty}=\min_{j\in[n]}\left\{\frac{\mathbf{M}(j,j)-\sum_{k\neq j}|\mathbf{M}(j,k)|}{\mathbf{M}(j,j)}\right\}>0$, which quantifies the strict row diagonal dominance of $\mathbf{M}$. The larger this quantity, the stronger the diagonal dominance and the faster the Neumann series converges. For strictly CDD $\mathbf{M}$, a similar quantity is $1-\left\|\mathbf{A}_{\mathbf{M}}^{\top}\mathbf{D}_{\mathbf{M}}^{-1}\right\|_{1}$. This hints us to consider the general quantity $1-\left\|\mathbf{D}_{\mathbf{M}}^{-1/q}\mathbf{A}_{\mathbf{M}}^{\top}\mathbf{D}_{\mathbf{M}}^{-1/p}\right\|_{p}$ for arbitrary $p\in[1,\infty]$. However, this quantity may be nonpositive for general RDD/CDD matrices $\mathbf{M}$, rendering it unsuitable for bounding the truncation error. To address this issue, we define the maximum $p$-norm gap by lazifying the involved operator and restricting it to $\operatorname{range}\left(\mathbf{I}-\mathbf{D}_{\mathbf{M}}^{-1/q}\mathbf{A}_{\mathbf{M}}^{\top}\mathbf{D}_{\mathbf{M}}^{-1/p}\right)$ (see Equation (5)), which guarantees positivity while remaining sufficient for bounding the truncation error.
Our algorithms estimate $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}=\frac{1}{2}\boldsymbol{t}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ by adapting three techniques for random-walk probability estimation on graphs: random-walk sampling [ST04, FRCS05], local push [ACL07, ABCHMT08], and the bidirectional method [LBG16].
@@
Based on local push methods, we formulate a Push primitive that estimates the vector $\boldsymbol{x}^{\ast}_{L}=\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ through deterministic local computation. Push maintains coordinate variables initialized to $\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ and iteratively applies the linear operator $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)$ to selected coordinates. Our analysis relies on invariant properties preserved by these operations, including an inequality variant that helps to handle negative entries. This approach yields closed-form accuracy guarantees for RDD systems and complexity bounds for special CDD systems. Combining these two aspects yield our result for special RCDD systems.
The bidirectional method combines random-walk sampling and local push from two directions. We adapt the BiPPR framework [LBG16], performing Push from $\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ and exploiting the invariant property to construct an unbiased estimator for $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$, which can be sampled via random walks from $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$ when $\mathbf{M}$ is RDD. By balancing the computational costs of both components, we achieve improved dependence on $L$ and $\varepsilon$, particularly for RCDD systems.